From a worksheet:
How long will it take $1750 to triple at 9.5% simple interest?

Answer by ankor@dixie-net.com(22740)

How long will it take $1750 to triple at 9.5% simple interest?
:
Simple interest formula: P*r*t = I
:
I assume that "Triple" means that you start with 1750, and end up with 5250
an additional $3500 added to the original 1750, $3500 is earned from interest
:
P = $1750; r = .095; I = $3500
:
1750 * .095 * t = 3500
:
166.25t = 3500
:
t = 3500/166.25
:
t = 21.05 years to triple your money
:
Check:
I = 1750 * .095 * 21.05
I = 3499.56 ~ $3500